OpenStudy (anonymous):

Find the derivative: x-2/x+2

OpenStudy (anonymous):

f(x)=x-2/x+2

OpenStudy (anonymous):

Just use the quotient rule............................................................ Dx[ f(x) / g(x)] = (f(x)*g'(x) - f'(x)*g(x)) / (g(x))^2

OpenStudy (anonymous):

ok I did that and I got 4/(x+2)^2